OpenStudy (anonymous):

the slope of the line is -8/7. find the point slope equation of the line. (4,4)

OpenStudy (compassionate):

y - 4 = -8/7(x - 4) ^ Just plug them in

OpenStudy (anonymous):

so its y - 4 = -8/7 (x-4) ?

OpenStudy (compassionate):

Yes